Ok so I was doing a problem where I needed to subtract 120*tan(40) - 60*tan(40)

So I used a calc and got
120*tan(40) = 100.69
60*tan(40) = 50.35
100.69-50.35 = 50.34

But wait, shouldn't the result be 50.35?
120tan(40) - 60tan(40) is the same as 60tan(40) right? and 60tan(40) = 50.35

I don't get why the discrepancy between the 50.34 and 50.35?

Found 3 solutions by ikleyn, MathTherapy, greenestamps:
Answer by ikleyn(52908)

.

120*tan(40°) - 60*tan40°) = (120-60)*tan(40°) = 60*tan(40°).             ANSWER


If you want to go further and to evaluate the last expression,
then find tan(40°) using your calculator, and substitute this value into the expression.



Answer by MathTherapy(10557)

Ok so I was doing a problem where I needed to subtract 120*tan(40) - 60*tan(40)

So I used a calc and got
120*tan(40) = 100.69
60*tan(40) = 50.35
100.69-50.35 = 50.34

But wait, shouldn't the result be 50.35?
120tan(40) - 60tan(40) is the same as 60tan(40) right? and 60tan(40) = 50.35

I don't get why the discrepancy between the 50.34 and 50.35?
 


Answer by greenestamps(13215)



The "discrepancy" was because you rounded your intermediate results.

In virtually all calculations, you want to keep as many digits of accuracy as you can until the very end; then round only once.

Your calculations:
120*tan(40) = 100.69
60*tan(40) = 50.35
100.69-50.35 = 50.34

Much better calculations:
120*tan(40) = 100.6919557
60*tan(40) = 50.34597787
100.6919557-50.34597787 = 50.34597787

That rounds to 50.35....

Of course, the easier calculation is

120*tan(40)-60*tan(40) = 60*tan(40) = 50.34597787

And then round that....